# Foundation Models in Medical Imaging - A Review and Outlook

### v3

@@ 3.2.4 Vision-Language Foundation Models
In addition to vision-only pathology FMs, research on multimodal learning from vision and language data in pathology has grown rapidly in recent years. As a result, several VLFMs, mostly pre-trained with cross-modal contrastive objectives, have been published [80, 81, 82, 83, 84, 85, 86, 87, 88, 89]. The majority of these works use tile-level vision encoders trained with image-text pairs sourced from, for example, pathology textbooks, research papers, YouTube videos, and medical Twitter.
Models such as MI-Zero and PLIP extend the CLIP framework to pathology. MI-Zero aligns patch-level embeddings from a vision encoder with text embeddings from a GPT-style transformer trained on pathology reports and PubMed abstracts, allowing zero-shot cancer subtyping without additional labels [81]. Similarly, PLIP, trained on data curated from Twitter, performs well on tasks like cancer classification and disease progression prediction [80]. Building on this, QUILTNET and CONCH further increase the scale at which VLFMs are trained, with QUILTNET fine-tuned on 1M image-text pairs (QUILT dataset) and excelling in cross-modal retrieval [82], and CONCH, trained on 1.17M pairs using the CoCa framework, achieving strong results in classification, retrieval, and tissue segmentation tasks [83].
Table 2: Vision-Language Pathology FMs. Data size shows amount of image-text pairs. Since these are collected from non-clinical sources such as PubMed, Twitter or Youtube, no distinction between WSI and patches is made.
@@
  • a
    iBOT is used for training their own image encoder
  • b
    Dataset is partly available. Weights are only available trained on model with open-source data, but not for model trained on full dataset.
More recently, generative approaches to VLFMs for pathology have also emerged. For instance, PathCLIP, a CLIP variant fine-tuned on 207K pathology image-caption pairs from PubMed, serves as the visual backbone for PathAsst, a multimodal generative model combining PathCLIP with the Vicuna-13b language model [85]. PathAsst enables advanced tasks such as visual question answering, multimodal dialogue, and image interpretation. Other instruction-tuned generative models, including PA-LLaVA [90], PathChat [91], and Quilt-LLaVA [92], show similar strong performance in visual question answering tasks.
@@ 3.2.5 Slide-level Foundation Models
The context modeled by tile-level pathology FMs is generally limited, resulting in the loss of contextual and spatial information beyond the image regions used for training. Although HIPT models a significantly larger context window, this method is still unable to capture global patterns at the slide level in a self-supervised fashion. Some methods have been introduced to overcome this. For example, the vision encoder LongVit, presented by [93], can directly learn slide representations end-to-end by leveraging LongNet and DINO. With its dilated attention and distributed algorithm, LongNet can process the extremely long token sequences resulting from gigapixel images, allowing LongVit to learn short- and long-range dependencies from millions of tiles.
@@
Recent models have combined patch- and slide-level encoders with vision-language alignment to enhance feature generalization and enable multimodal tasks. For example, Prov-GigaPath [86] uses a vision transformer (DINOv2) for tile-level features and a LongNet model to capture slide-level features, with CLIP-based vision-language alignment improving zero-shot performance in cancer subtyping and gene mutation prediction. This model outperforms patch-level FMs like REMEDIS and UNI in cancer detection and tumor origin prediction. Similarly, PRISM [84] aggregates tile-level embeddings from Virchow to generate slide-level embeddings for zero-shot cancer subtyping, detection, and clinical report generation. CHIEF [94] builds on this framework by pretraining a patch-level encoder, CTransPath, and combining it with weakly supervised pretraining on 60K WSIs. It also uses CLIP-based text prompts to encode anatomical site information, improving generalization for cancer detection and tumor origin prediction. TITAN [95] integrates histology knowledge distillation with contrastive vision-language learning to align regions of interest with captions and pathology reports, excelling in cancer subtyping, rare disease classification, and multimodal tasks like report generation.
@@
Other models enhance slide-level representation learning by integrating additional biological context or modalities. KEEP [96] incorporates a disease knowledge graph to enhance performance in rare cancer detection and subtyping. mSTAR [97] integrates WSIs, pathology reports, and RNA-Seq data, significantly improving unimodal and multimodal tasks, particularly genomic information use. THREADS [98] introduces molecular supervision by aligning whole-slide image features with matched transcriptomic and genomic profiles, enabling representations that reflect both morphological and molecular characteristics. Despite having a smaller slide encoder than those of PRISM and GigaPath, THREADS achieves superior performance on oncology tasks, including rare event prediction, and outperforms lightweight models like CHIEF.
@@
Finally, COBRA [99] represents an ensemble method that aggregates patch embeddings from multiple foundation models in feature space using a contrastive pre-training strategy. Tested on FMs CTransPath, UNI, Virchow2, and H-Optimus-0, COBRA is flexible and applicable to any ensemble, demonstrating strong performance despite being pre-trained on only 3048 WSIs.
@@
Table 3: WSI-Level Pathology Foundation Models. The dataset size is given in number of patches and the number of WSIs in parentheses.
Model Type Backbone SSL Pre-training Data Availability
Size a Source Weights Data
Giga-SSL VFM ResNet-18 SimCLR - (12K) TCGA
LongViT VFM LongNet DINO 1M (10K) TCGA
PRISM VLFM Virchow + BioGPT CoCa - (590K) Proprietary
Prov-GigaPath VLFMa ViT-G + PubMedBERT DINOv2 + ITC 1.3B (170K) Prov-Path
mSTAR VLFM ViT-L mSTAR - (26K) TCGA
KEEP VLFM Uni + PubMedBERT Knowledge-enhanced ITC - (143K) OpenPath, Quilt-1M
CHIEF VLFMa CTransPath + CLIP text encoder SRCL (v)+ ITC (vl) 15M (61K) Proprietary
TITAN VLFMa ViT-B + CONCHv1.5 iBOT (v) + CoCa (vl) - (60K) GTEx
COBRA VLFM ABMIL + Mamba-2 COBRA - (3K) mTCGA
THREADS VFM CONCHV1.5 (patch) + cGPT (gene) Contrastive -(47K) MBTG-47K b ✓/✗c
  • a
    Authors train both an image encoder as well as a VL-alignment module
  • b
    Planned to be released
  • c
@@
    Partially available: TCGA and GTEx are open-source, but pretraining data from BWH and MGH are proprietary.
@@ 3.2.6 Applying SAM in Pathology
Besides the pathology VFMs and VLFMs discussed above, some other works investigated the use of SAM for segmentation tasks in the pathology domain. This was done by either directly applying an out-of-the-box SAM model trained on natural images [100, 101, 102, 103] or by in-domain adaptation and fine-tuning of SAM on histopathology images [104, 105, 106, 107, 108, 109].
@@ 3.3 Concluding Remarks
Research on SSL and FM development in pathology has made great progress in recent years and has proven fruitful in several digital pathology applications. Early models like CTransPath laid the groundwork for domain-specific adaptations, while later efforts such as Virchow introduced scalability and multi-task capabilities at the slide level. The focus has shifted toward scaling both data and model size, as seen, for example, in the performance of UNI and Phikon, which adopt broader vision foundation model frameworks with specific adaptations for cancer and pathology tasks. Vision foundation models have been at the forefront of these efforts, with notable success in improving performance across various pathology tasks. However, there have also been works using vision and language, such as using vision-language alignment to enhance whole-slide FMs and developing visual question answering FMs for pathology.
Despite these advancements, challenges remain. Scaling datasets and model sizes has brought notable performance improvements. Still, results suggest a plateau in gains, emphasizing the need for better data curation, particularly to represent rare pathologies and diverse staining protocols. Additionally, foundational algorithms optimized for natural images may require further adaptation to fully exploit the unique characteristics of histopathology data, including high resolution and spatial redundancy. Future work will require balancing scale with task specificity while addressing technical hurdles such as standardizing data across institutions.
@@ 4 Radiology
Radiology data is obtained through various imaging techniques, such as plain radiography (X-ray) and computed tomography (CT), magnetic resonance imaging (MRI), ultrasound (US), and nuclear medicine imaging (such as PET scans) ([110]. Each modality provides data in different formats and dimensions: X-rays and CT scans use the same ionizing radiation to produce two-dimensional images and three-dimensional volumetric images, respectively, enabling for example detection of fractures, infections, tumors, and blood clots. MRI uses magnetic fields and radiofrequency waves to generate detailed 3D images of soft tissues, ideal for examining the brain, spine, muscles, and joints. Ultrasound uses high-frequency sound waves for 2D and 3D imaging, commonly used in prenatal check-ups and cardiac assessments. Nuclear medicine employs radioactive tracers to evaluate organ function, which is particularly useful for diagnosing cancer and cardiovascular and brain diseases. X-rays, CT, MRI, and US scans are the most frequently performed imaging tests in clinical practice.
@@ 4.1 Early Advances in Radiology Models
Unlike pathology, where the development of foundation models has seen significant progress, radiology has seen relatively fewer developments, with less foundation models emerging. Early radiology models primarily focused on leveraging transformer architectures for tasks like report generation and segmentation using CNN-transformer hybrids such as [111], [112], and [113]. However, these models did not utilize self-supervised learning, which limited their scalability. As research progressed, the focus shifted towards developing SSL techniques specialized for radiological imaging, particularly contrastive vision-language learning. Early work by [114] improved feature learning by leveraging part-whole relationships. [115] introduced ConVIRT to align medical images with textual descriptions, followed by GLoRIA [116], which refined this alignment at the sub-region level in radiology reports. Later, [117] focused on chest X-rays, enhancing semantic modeling through contrastive learning, which was further improved by Bio-ViL-T [30] with the addition of temporal dynamics, although the scale of these methods remained limited compared to modern foundation models.
@@ 4.2.1 Generalist Vision-Language Foundation Models
As foundation models became more popular, within the medical domain, a category of models emerged that emphasizes scale by incorporating a wide variety of medical imaging modalities. These generalist vision-language models utilize extensive paired datasets to address various medical imaging modalities, including radiology, pathology and ophthalmology. Models such as PubMed-CLIP [118], PMC-CLIP [119], BiomedCLIP [87], and MEDVInT [120] utilized extensive datasets from PubMed to perform multimodal tasks like disease classification, report generation, image retrieval, and medical question answering (MedVQA). Models such as LLaVA-Med [121], which finetuned LLaVA [47], and Qilin-Med-VL [122], which used curriculum learning and instruction tuning, specialize in MedVQA tasks. Similarly, Med-Flamingo [123], built further on the trained Flamingo architecture [28] for general medical VQA. Although evaluated on radiological tasks such as X-ray, these models were not specifically tailored for radiological imaging, focusing more broadly on multimodal medical datasets. Additionally, these models did not utilize 3D data, instead relying on 2D slices of 3D modalities such as CT and MRI.
Table 4: Foundation Models for Multi-Domain Radiology
Model Type Imaging Domain Backbone SSL Objective Data Size Availability

### v4

@@ 3.2.4 Vision-Language Foundation Models
In addition to vision-only pathology FMs, research on multimodal learning from vision and language data in pathology has grown rapidly in recent years. As a result, several VLFMs, mostly pre-trained with cross-modal contrastive objectives, have been published [huang2023visual, lu2023visual, ikezogwo2024quilt, lu2024visual, shaikovski2024prism, sun2024pathasst, xu2024whole, zhang_biomedclip_2024, qu2024rise, zhang2023text]. The majority of these works use tile-level vision encoders trained with image-text pairs sourced from, for example, pathology textbooks, research papers, YouTube videos, and medical Twitter.
Models such as MI-Zero and PLIP extend the CLIP framework to pathology. MI-Zero aligns patch-level embeddings from a vision encoder with text embeddings from a GPT-style transformer trained on pathology reports and PubMed abstracts, allowing zero-shot cancer subtyping without additional labels [lu2023visual]. Similarly, PLIP, trained on data curated from Twitter, performs well on tasks like cancer classification and disease progression prediction [huang2023visual]. Building on this, QUILTNET and CONCH further increase the scale at which VLFMs are trained, with QUILTNET fine-tuned on 1M image-text pairs (QUILT dataset) and excelling in cross-modal retrieval [ikezogwo2024quilt], and CONCH, trained on 1.17M pairs using the CoCa framework, achieving strong results in classification, retrieval, and tissue segmentation tasks [lu2024visual].
Table 2: Vision-Language Pathology Foundation Models. Data size shows amount of image-text pairs. Since these are collected from non-clinical sources such as PubMed, Twitter or Youtube, no distinction between WSI and patches is made.
@@
a iBOT is used for training their own image encoder.  b Dataset is partly available. Weights are only available trained on model with open-source data, but not for model trained on full dataset.
More recently, generative approaches to VLFMs for pathology have also emerged. For instance, PathCLIP, a CLIP variant fine-tuned on 207K pathology image-caption pairs from PubMed, serves as the visual backbone for PathAsst, a multimodal generative model combining PathCLIP with the Vicuna-13b language model [sun2024pathasst]. PathAsst enables advanced tasks such as visual question answering, multimodal dialogue, and image interpretation. Other instruction-tuned generative models, including PA-LLaVA [dai2024pa], PathChat [lu2024multimodal], and Quilt-LLaVA [seyfioglu2024quilt], show similar strong performance in visual question answering tasks.
@@ 3.2.5 Slide-level Foundation Models
The tile-level pathology FMs discussed thus far are scalable and computationally efficient, but their reliance on local tiles means they lose global spatial context, limiting performance on tasks that require holistic assessment of tissue architecture. Some hierarchical tile-level models, such as HIPT, partially address this limitation by aggregating multi-scale information, extending the effective context window. However, HIPT still operates primarily at the patch level and cannot capture truly global slide-level patterns in a single self-supervised model.
Slide-level approaches address this by directly modeling entire WSIs, aiming to capture both local and global dependencies. One such method is the vision encoder LongVit [wang2023image], which leverages DINO and LongNet [ding2023longnet] to process the extremely long token sequences produced by gigapixel images, enabling end-to-end slide-level representation learning. While LongVit can model a larger context than conventional tile-level encoders, slide-level self-supervised modeling remains computationally demanding and less widely benchmarked than tile-level approaches.
@@
Recent models have combined patch- and slide-level encoders with vision-language alignment to enhance feature generalization and enable multimodal tasks. For example, Prov-GigaPath [xu2024whole] uses a vision transformer (DINOv2) for tile-level features and a LongNet model to capture slide-level features, with CLIP-based vision-language alignment improving zero-shot performance in cancer subtyping and gene mutation prediction. This model outperforms patch-level FMs like REMEDIS and UNI in cancer detection and tumor origin prediction. Similarly, PRISM [shaikovski2024prism] aggregates tile-level embeddings from Virchow to generate slide-level embeddings for zero-shot cancer subtyping, detection, and clinical report generation. CHIEF [wang2024pathology] builds on this framework by pretraining a patch-level encoder, CTransPath, and combining it with weakly supervised pretraining on 60K WSIs. It also uses CLIP-based text prompts to encode anatomical site information, improving generalization for cancer detection and tumor origin prediction. TITAN [ding2024multimodal] integrates histology knowledge distillation with contrastive vision-language learning to align regions of interest with captions and pathology reports, excelling in cancer subtyping, rare disease classification, and multimodal tasks like report generation.
@@
Other models enhance slide-level representation learning by integrating additional biological context or modalities. KEEP [zhou2024knowledgeenhancedpathologyvisionlanguagefoundation] incorporates a disease knowledge graph to enhance performance in rare cancer detection and subtyping. mSTAR [xu2024multimodal] integrates WSIs, pathology reports, and RNA-Seq data, significantly improving unimodal and multimodal tasks, particularly genomic information use. THREADS [vaidya2025molecular] introduces molecular supervision by aligning whole-slide image features with matched transcriptomic and genomic profiles, enabling representations that reflect both morphological and molecular characteristics. Despite having a smaller slide encoder than those of PRISM and GigaPath, THREADS achieves superior performance on oncology tasks, including rare event prediction, and outperforms lightweight models like CHIEF.
@@
Finally, COBRA [lenz2024unsupervised] represents an ensemble method that aggregates patch embeddings from multiple foundation models in feature space using a contrastive pre-training strategy. Tested on FMs CTransPath, UNI, Virchow2, and H-Optimus-0, COBRA is flexible and applicable to any ensemble, demonstrating strong performance despite being pre-trained on only 3048 WSIs.
@@
Table 3: WSI-Level Pathology Foundation Models. Type refers to whether the FM is a vision model (VM) or a vision-language model (VLM). Backbone and SSL objective show how the model was pretrained. The pre-training dataset size is given in number of patches and the number of WSIs in parentheses. Availability shows whether the weights and pretraining data are available online.
Model Type Backbone SSL Pre-training Data Availability Size Source Weights Data Giga-SSL VM ResNet-18 SimCLR - (12K) TCGA LongViT VM LongNet DINO 1M (10K) TCGA PRISM VLM Virchow + BioGPT CoCa - (590K) Proprietary Prov-GigaPath VLMa ViT-G + PubMedBERT DINOv2 + ITC 1.3B (170K) Prov-Path mSTAR VLM ViT-L mSTAR - (26K) TCGA KEEP VLM Uni + PubMedBERT Knowledge-enhanced ITC - (143K) OpenPath, Quilt-1M CHIEF VLMa CTransPath + CLIP text encoder SRCL (v)+ ITC (vl) 15M (61K) Proprietary TITAN VLMa ViT-B + CONCHv1.5 iBOT (v) + CoCa (vl) - (60K) GTEx COBRA VLM ABMIL + Mamba-2 COBRA - (3K) mTCGA THREADS VM CONCHV1.5 (patch) + cGPT (gene) Contrastive - (47K) MBTG-47K b ✓/✗c
@@
a Authors train both an image encoder as well as a VL-alignment module. b Planned to be released. c Partially available: TCGA and GTEx are open-source, but pretraining data from BWH and MGH are proprietary.
@@ 3.2.6 Applying SAM in Pathology
The Segment Anything Model (SAM) is a vision foundation model originally designed for universal image segmentation, trained on over a billion masks from natural images [kirillov2023segment]. Its flexibility in generating segmentation masks from simple prompts makes it appealing for medical applications, even though it does not fully align with the self-supervised pretraining paradigm central to our review.
Consequently, SAM has been explored either via direct zero-shot use on histopathology images [chauveau2023segment, deng2023segment, zhang2023input, zhu2023samms] or through in-domain fine-tuning on pathology-specific datasets [ranem2023exploring, zhang2023sam, chen2024sam, cui2024all, su2024adapting, zhang2024sam2]. These works show that SAM can deliver promptable segmentation without extensive annotation. However, its performance remains limited in addressing histopathological complexities, varying scales, and dense cellular structures, making SAM most useful as a transfer baseline rather than a domain-specific foundation model.
@@ 3.3 Concluding Remarks
Research on SSL and FM development in pathology has made great progress in recent years and has proven fruitful in several digital pathology applications. Early models like CTransPath laid the groundwork for domain-specific adaptations, while later efforts such as Virchow introduced scalability and multi-task capabilities at the slide level. The focus has shifted toward scaling both data and model size, as seen, for example, in the performance of UNI and Phikon, which adopt broader VFM frameworks with specific adaptations for cancer and pathology tasks. Vision foundation models have been at the forefront of these efforts, with notable success in improving performance across various pathology tasks. However, there have also been works using vision and language, such as using vision-language alignment to enhance whole-slide FMs and developing visual question answering FMs for pathology.
Despite these advancements, challenges remain. Scaling datasets and model sizes has brought notable performance improvements. Still, results suggest a plateau in gains, emphasizing the need for better data curation, particularly to represent rare pathologies and diverse staining protocols. Additionally, foundational algorithms optimized for natural images may require further adaptation to fully exploit the unique characteristics of histopathology data, including high resolution and spatial redundancy. Tile-level FMs have so far proven more scalable and robust, as they can be trained efficiently on individual patches, require lower computational resources than slide-level models, and have demonstrated strong performance across diverse datasets and benchmark tasks. In contrast, slide-level FMs promise richer contextual understanding by modeling entire WSIs, but remain limited by their computational demands, smaller training cohorts, and fewer established benchmarks. Future work will require balancing scale with task specificity while addressing technical hurdles such as standardizing data across institutions.
@@ 4 Radiology
Radiology data is obtained through various imaging techniques, such as plain radiography (X-ray) and computed tomography (CT), magnetic resonance imaging (MRI), ultrasound (US), and nuclear medicine imaging (such as PET scans) [islam2023introduction]. Each modality provides data in different formats and dimensions: X-rays and CT scans use the same ionizing radiation to produce two-dimensional images and three-dimensional volumetric images, respectively, enabling for example detection of fractures, infections, tumors, and blood clots. MRI uses magnetic fields and radiofrequency waves to generate detailed 3D images of soft tissues, ideal for examining the brain, spine, muscles, and joints. Ultrasound uses high-frequency sound waves for 2D and 3D imaging, commonly used in prenatal check-ups and cardiac assessments. Nuclear medicine employs radioactive tracers to evaluate organ function, which is particularly useful for diagnosing cancer and cardiovascular and brain diseases. X-rays, CT, MRI, and US scans are the most frequently performed imaging tests in clinical practice.
@@ 4.1 Early Advances in Radiology Models
Unlike pathology, where the development of foundation models has seen significant progress, radiology has seen relatively fewer developments, with less foundation models emerging. Early radiology models primarily focused on leveraging transformer architectures for tasks like report generation and segmentation using CNN-transformer hybrids such as [xie_cotr_2021], [hatamizadeh_unetr_2021], and [chen_transunet_2021]. However, these models did not utilize self-supervised learning, which limited their scalability. As research progressed, the focus shifted towards developing SSL techniques specialized for radiological imaging, particularly contrastive vision-language learning. Early work by [feng_parts2whole_2020] improved feature learning by leveraging part-whole relationships. [zhang2020contrastive] introduced ConVIRT to align medical images with textual descriptions, followed by GLoRIA [huang_gloria_2021], which refined this alignment at the sub-region level in radiology reports. Later, [boecking2022making] focused on chest X-rays, enhancing semantic modeling through contrastive learning, which was further improved by Bio-ViL-T [bannur2023learning] with the addition of temporal dynamics, although the scale of these methods remained limited compared to modern foundation models.
@@ 4.2.1 Generalist Vision-Language Foundation Models
As foundation models became more popular, within the medical domain, a category of models emerged that emphasizes scale by incorporating a wide variety of medical imaging modalities. These generalist vision-language models utilize extensive paired datasets to address various medical imaging modalities, including radiology, pathology and ophthalmology. Models such as PubMed-CLIP [eslami_does_2021], PMC-CLIP [lin2023pmc], BiomedCLIP [zhang_biomedclip_2024], and MEDVInT [zhang_pmc-vqa_2023] utilized extensive datasets from PubMed to perform multimodal tasks like disease classification, report generation, image retrieval, and medical question answering (MedVQA). Models such as LLaVA-Med [li_llava-med_2023], which finetuned LLaVA [liu2023visual], and Qilin-Med-VL [liu_qilin-med-vl_2023], which used curriculum learning and instruction tuning, specialize in MedVQA tasks. Similarly, Med-Flamingo [moor_med-flamingo_2023], built further on the trained Flamingo architecture [alayrac2022flamingo] for general medical VQA. Although evaluated on radiological tasks such as X-ray, these models were not specifically tailored for radiological imaging, focusing more broadly on multimodal medical datasets. Additionally, these models did not utilize 3D data, instead relying on 2D slices of 3D modalities such as CT and MRI.
Table 4: Foundation Models for Multi-Domain Radiology. Type refers to whether the FM is a vision model (VM) or a vision-language model (VLM). Backbone and SSL objective show how the model was pretrained. Imaging domain shows which medical image types were used to pre-train the model. Dataset sizes are given as the number of 2D images used for pre-training, unless specified otherwise. Availability shows whether the weights and pretraining data are available online.
Model Type Imaging Domain Backbone SSL Objective Data Size Availability Weights Data PubMedCLIP VLM 2D Radiologyc ResNet50 + ViT-B/32 ITC 80K UniMiSS VM CT(3D), X-Ray MiT MIM 114Kd BiomedCLIP VLM Multimodalb ViT-B/16 + PubMedBERT ITC 15M PMC-CLIP VLM 2D Radiologyc ResNet50 + PubMedBERT ITC + MLM 1.65M MEDVInT VLM Multimodalb ResNet50 + Transformer MLM 227K LLaVA-Med VLM Multimodalb LLaVA ITC 600K LVM-Med VM CT(2D), MRI, X-ray, US ResNet-50 + ViT Graph Matching 1.3M Med-Flamingo VLM Multimodalb ViT/L-14 + LLaMA-7B ITC 1.6M RadFM VLM X-Ray, CT, MRI (3D), US, PET 3D ViT + MedLLaMA-13B Generative ITC 16Me Qilin-Med-VL VLM Multimodalb ViT/L-14 + LLaMA-13B ITC 580K SAT VLM Multimodalb 3D U-Net + BioBERT ITC 302K VISION-MAE VM CT, MRI, PET, X-Ray, US Swin-T MAE 2.5M RadCLIP VLM CT (2D+3D), X-Ray, MRI ViT-L/14 Contrastive 1.2M
(a) X-Ray, US, MRI, PET, fluoroscopy, mammography, angiography. (b) Multimodal = multiple radiology + pathology modalities. (c) Includes CT, MRI, PET, X-Ray, US, fMRI, etc. (d) Dataset includes 109K 2D images + 5K 3D volumes. (e) Includes 15.5M 2D and 500K 3D image-text pairs.
4.2.2 Chest X-Ray Vision-Language Foundation Models
One subdomain within radiology that has seen relatively more FM related developments than other modalities is Chest X-ray (CXR). CXR lends itself well to foundation model research due to its widespread use in clinical settings, the availability of large and well-established datasets such as [irvin2019chexpert, johnson2019mimic], and its relatively reduced computational complexity compared to more intricate 3D modalities like CT and MRI. However, unlike the earlier mentioned generalist models that can scale by incorporating multiple imaging modalities, gathering large, paired CXR-specific datasets remains challenging. As a result, some models like MedCLIP [wang_medclip_2022] have addressed this by leveraging unpaired data with semantic matching, enabling the use of more diverse datasets. Other models focused on enhancing the quality of existing paired data to maximize the learning potential and improve feature extraction. For example, CheXzero [tiu_expert-level_2022] employed a contrastive learning approach using only the concise "impressions" section of radiology reports, leading to better feature extraction and zero-shot disease classification and CXR-CLIP [you_cxr-clip_2023] built on this by creating high-quality image-text pairs with radiologist-designed prompts, further improving feature retrieval and interpretability. KAD [zhang_knowledge-enhanced_2023] incorporated a medical knowledge graph into its training process, allowing for deeper reasoning and improved visual representation learning, although its reliance on single-source data limited its adaptability. To overcome this, UniChest [dai_unichest_2024] introduced a conquer-and-divide framework that systematically captured shared patterns across diverse datasets while isolating source-specific variations, resulting in better generalization across CXR benchmarks. Similarly, MedKLIP [wu2023medklip] focused on fine-grained alignment between images and text to enhance interpretability and disease classification, while MaCo [huang_enhancing_2023] introduced a masked contrastive approach, refining image-text alignment and further improving zero-shot classification performance.
In addition to knowledge-enhancing models, others like ELIXR [xu2023elixr] and MAIRA-1 [hyland2023maira] focused on efficient adaptation by using specialized adapters to align visual encoders with large language models, reducing the need for full retraining. A more recent adaptation of MAIRA-1, MAIRA-2, combines two pre-trained image- and language encoders and adapts them for report generation from chest X-ray images [bannur2024maira]. More recently, Google’s CheXAgent [chen_chexagent_2024] marked a shift toward scaling with large datasets. By leveraging the CheXinstruct dataset, which includes 6 million instruction-image-answer triplets, CheXAgent focused on large-scale instruction-tuned learning and demonstrated superior performance across multiple clinically relevant tasks, highlighting the potential of large-scale, instruction-based approaches.
Table 5: Foundation Models specialized for Chest X-Ray. Type refers to whether the FM is a vision model (VM) or a vision-language model (VLM). Backbone and SSL objective show how the model was pretrained. Data sizes are given as the number of 2D images used for pre-training. Availability shows whether the weights and pretraining data are available online.
Model Type Backbone SSL Objective Data Size Availability